The sides of a square are increasing at a rate of 10 cm/sec. How fast is the area enclosed by the square increasing when the area is 150 cm^2.



1
Expert's answer
2021-10-11T06:06:55-0400

The area of a square:

"S=a^2"

where a is side of a square.


"\frac{dS}{dt}=2a\frac{da}{dt}"


"a=\sqrt{S}=\sqrt{150}=5\sqrt{6}" cm


"\frac{da}{dt}=10" cm/s


"\frac{dS}{dt}=2\cdot 5\sqrt{6}\cdot 10=100\sqrt{6}" cm2/s